OpenStudy (anonymous):

integrate: (x-1)/(x^2 + 1) I recognize the denominator as arctan but not sure what that gets me?

OpenStudy (anonymous):

1/2 ln(x^2+1) - arctan(x)

OpenStudy (mimi_x3):

You can break it up: \[\large \int\limits\frac{x-1}{x^{2}+1} dx = \int\limits\frac{x}{x^{2}+1} dx -\int\limits\frac{1}{x^{2}+1} dx\] The integrate it one by one.

OpenStudy (mimi_x3):

The first one is a u sub & the next one is a trig inverse.

OpenStudy (anonymous):

I was trying to make it way harder than that lol...dangit

OpenStudy (mimi_x3):

Do you want to do it a hard way? lol